# Using Dynamic Safety Margins as Control Barrier Functions

### v3

@@ 2.3 Dynamic Safety Margins
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$ is hereafter referred to as a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
@@
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{v}}]$.
@@
We also define the state-reference constraint set
$\tilde{\mathcal{X}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},\boldsymbol{\mathbf{v}}\in\mathcal{V}\}.$ (21)
The following DSM definition satisfies the one proposed in [23, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
@@
[…]
@@
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\end{bmatrix},$ (23)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big{(}\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big{)}$. In the ERG literature, $\rho$ is called a navigation field [23, Definition 2]. System (23) ensures that, if $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$, then $\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big{)}\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [23]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [30, 23, 31, 32, 33, 34]. The following section shows how these tools from the ERG literature can be used to construct valid CBFs.
@@ 3.1 DSMs are CBFs
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown that the set-valued maps $\mathcal{C}_{i}$ are both upper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29] that each $\hat{\alpha}_{i}$ is continuous on $[0,\infty)$. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, and it follows by (22c) that
@@
There always exists $\alpha_{i}\in\mathscr{K}_{\infty}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}_{\infty}$. Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, it follows that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. With this,
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$. Then, there exists $\alpha\in\mathscr{K}_{\infty}$ such that $\Delta$ satisfies (9). Specifically, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Moreover, if 00 is a regular value of each $\Delta_{i}$, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
Proof 3.9.
The first statement was already shown in the proof of Theorem 3.6. The control invariance of $\tilde{\mathcal{C}}$ is an immediate consequence of Theorem 2 when $n_{c}=1$. For the general case $n_{c}>1$, we refer the reader to [28, Proposition 4.3.7].
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$ and let $\alpha\in\mathscr{K}_{\infty}$ be such that (9) holds. If $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ is such that
@@ 3.2 DSM-CBF Optimization-based Policy
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}_{\infty}$ such that the optimization-based policy $\beta:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
@@
is feasible for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$. Moreover, if $\beta$ is locally Lipschitz continuous and 00 is a regular value of each $\Delta_{i}$, then the trajectory of the augmented system (24) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$.
Proof 3.13.
@@
The proposed approach can be interpreted as a closed-form variant of the backup CBF [21]. The approach assigns a safe control input $\boldsymbol{\mathbf{u}}$ while simultaneously generating a safe reference $\boldsymbol{\mathbf{v}}$ for the backup controller $\pi$.
Remark 3.18.
@@ 3.3 Existence and Uniqueness of Solutions
    (Slater CQ) There exists $\alpha\in\mathscr{K}_{\infty}$ such that for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$;
@@
    (Constant Rank CQ) There exists a neighborhood of every point in $\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{I}\subset\{1,\ldots,n_{c}\}$, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix},i\in\mathcal{I}\right\}$ remains constant;
@@ 4 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [23, 30, 31, 32] can be used within the context of this paper.
@@
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}_{\infty}$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\boldsymbol{\mathbf{v}})$, denoted $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$, wherein
@@
Additionally, we require $V$ has non-vanishing gradient near the equilibrium points. Formally, we have the following assumption.
@@
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}_{\infty}$ can be used for the DSM-CBF policy $\beta$ given in (31). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
@@
If $\Delta$ is a Lyapunov-based DSM as in (37), then $\forall\alpha\in\mathscr{K}_{\infty}$, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$.
Proof 4.28.
Let $\alpha\in\mathscr{K}_{\infty}$ and $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
@@
The proposed method effectively transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [30, 31, 32] and numerical [34] tools.

### v4

@@ 2.3 Dynamic Safety Margins
Assumption 1 is standard in the reference governor literature: it applies to any system that admits a connected path of stabilizable equilibrium points. As noted in [4], this encompasses a wide range of meaningful systems and applications.
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$ is hereafter called a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
@@
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $\boldsymbol{\mathbf{x}};~\boldsymbol{\mathbf{v}}]$.
@@
Lemma 2.6.
The state-reference constraint set
$\tilde{\mathcal{X}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},~\boldsymbol{\mathbf{v}}\in\mathcal{V}\},$ (21)
is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$.
Proof 2.7.
Let $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\tilde{\mathcal{X}}$ be a sequence that converges to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{X}$ and $\{\boldsymbol{\mathbf{v}}_{n}\}\subset\mathcal{V}$, and $\mathcal{X}$, $\mathcal{V}$ are closed, it follows that $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Define the sequence $\{\pi(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\mathcal{U}$ and note that by continuity of $\pi$, it converges to $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ because $\mathcal{U}$ is closed. With this, we have that $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ and it follows that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{X}}$.
The following DSM definition satisfies the one proposed in [24, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
@@
$\tilde{\mathcal{C}}~\text{is compact in}~\mathbb{R}^{n}\times\mathbb{R}^{l},$ (22b)
@@
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix}\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}}),$ (23)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big(\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big)$. In the ERG literature, $\rho$ is called a navigation field [24, Definition 2]. System (23) ensures that, if $\big(\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big)\in\tilde{\mathcal{C}}$, then $\big(\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big)\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [24]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [32, 24, 33, 34, 35, 36]. The following section shows how these tools from the ERG literature can be used to construct valid CBFs.
@@ 3.1 DSMs are CBFs
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown that the set-valued maps $\tilde{\mathcal{C}}_{i}$ are both upper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [31] that each $\hat{\alpha}_{i}$ is continuous on $0,\infty)$. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, and it follows by (22c) that
@@
There always exists $\alpha_{i}\in\mathscr{K}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}$. Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, it follows that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. With this,
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$. Then, there exists $\alpha\in\mathscr{K}$ such that $\Delta$ satisfies (9). Specifically, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\big(\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Moreover, if 0 is a regular value of each $\Delta_{i}$, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
Proof 3.12.
The first statement was already shown in the proof of Theorem 3.9. The control invariance of $\tilde{\mathcal{C}}$ is an immediate consequence of Theorem 2 when $n_{c}=1$. For the general case $n_{c}>1$, we refer the reader to [30, Proposition 4.3.7].
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$ and let $\alpha\in\mathscr{K}$ be such that (9) holds. If $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ is such that
@@ 3.2 DSM-CBF Optimization-based Policy
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}$ such that the optimization-based policy $\beta:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
@@
is feasible for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$. Moreover, if $\beta$ is locally Lipschitz continuous and 0 is a regular value of each $\Delta_{i}$, then the trajectory of the augmented system (24) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big(\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big)\in\tilde{\mathcal{C}}$.
Proof 3.16.
@@ 3.3 Existence and Uniqueness of Solutions
    (Slater CQ) There exists $\alpha\in\mathscr{K}$ such that for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$;
@@
    (Constant Rank CQ) There exists a neighborhood of every pair $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{I}\subset\mathcal{J}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, where $\mathcal{J}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is the set of active constraints, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix},i\in\mathcal{I}\right\}$ remains constant;
@@ 4 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [24, 32, 33, 34] can be used within the context of this paper. For trajectory-based DSMs, refer to [41] instead, which can be regarded as a generalization of backup CBFs [22] when using dynamic backup policies.
@@
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\boldsymbol{\mathbf{v}})$, denoted $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$, wherein
@@
Additionally, we require $V$ has non-vanishing gradient near the equilibrium points with the following assumption.
@@
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}$ can be used for the DSM-CBF policy $\beta$ given in (31). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
@@
If $\Delta$ is a Lyapunov-based DSM as in (37), then $\forall\alpha\in\mathscr{K}$, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$.
Proof 4.30.
Let $\alpha\in\mathscr{K}$ and $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
@@
The proposed method effectively transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [32, 33, 34] and numerical [36] tools. It is also worth noting that one can always linearize the system around the current reference $v$ and use the Lyapunov function of the linearized system.